Please properly answer the questions listed below.

Also, please (TYPE) everything out.


3- Suppose that the government is running a balanced budget and the value of purchases made by the government is 200. The consumption function is C = 200 + 0.6 Yd and planned investment is 100.
b) If the level of aggregate output is 1250, calculate:
- Disposal income
- Aggregate consumption
- Aggregate Saving
- Planned aggregate expenditure
- Unplanned inventory change
Y = C + S + T
Yd = Y – T
Yd = C + S
AE = C + I + G
Unplanned change in inventory: Y – (C + I + G)
4- Knowing that Y = C + S + T and AE = C + I + G, show that in equilibrium the sum of savings and taxes is equal to the sum of planned investment and government purchases.
5- Suppose a consumption function of an economy is C = 500 + 0.8 Yd ,
a) Calculate the government spending multiplier.
b) Calculate the tax multiplier.
c) Calculate the balanced-budget multiplier.
d) If the government increases it purchases for 200, how much the equilibrium outcome will change?
e) If the government decreases taxes for 200, how much the equilibrium outcome will change?
f) If the government increases it purchases for 200 and keep the budget balanced, how much the equilibrium outcome will change?
Answer #1

3) A balanced budget means government spending equals to tax. So G=200=T

Disposal income yd=Y-T=1250-200=1050

Aggregate consumption C=200+0.6Yd = 200+(0.6*1050)= 830

Aggregate savings= Yd-C= 1050-830= 220

Planned aggregate expenditure AE=C+I+G=830+100+200= 1130

Unplanned inventory changes Y-(C+I+G)=1250-(830+100+200)=120